OpenStudy (anonymous):

¿Cuál es la solución a la ecuación: 2x(x-7)-2x2=14 ?

OpenStudy (ksaimouli):

x=-1

OpenStudy (ksaimouli):

2x2−14x−2x2=14 −14x=14 x=14÷−14 x=−1

OpenStudy (anonymous):

-1